OpenStudy (anonymous):

Simplify: (w^3-w^2y)/(wz-w) /((w-y)/(z-1))^2

OpenStudy (espex):

Is the problem \[\large{\frac{\frac{w^3-w^2y}{wz-w}}{(\frac{w-y}{z-1})^2}}\]

OpenStudy (anonymous):

\[\left(\begin{matrix}w^3-w^2y \\ wz-w\end{matrix}\right)\div \left(\begin{matrix}w-y \\ z-1\end{matrix}\right)^2\]

OpenStudy (anonymous):

=[w^2*(w-y)/w*(z-1)]/[(w-y)/(z-1)]^2 =[w*(w-y)/(z-1)]/[(w-y)/(z-1)]^2 =w*(z-1)/(w-y)